OpenStudy (anonymous):

•Need help with 3 questions •Need 2 Questions checked •Will Medal & Fan!

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Ok thanks, now what about the other 3? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

give it a shot and tell me what you get

jimthompson5910 (jim_thompson5910):

or tell me how far you get with it

OpenStudy (anonymous):

ooh that one about the linear function was tricky

OpenStudy (anonymous):

i'd have said "sometimes true" also because \(f(x)=2\) does not have an inverse for example but if you read carefully it does not say "inverse function" it says "inverse RELATION"

OpenStudy (anonymous):

RIght, these are kinda tricky.

jimthompson5910 (jim_thompson5910):

yeah the `relation` keyword makes me think it's always true because `y = 2` will have an inverse of `x = 2` which isn't a function, but it's a relation

OpenStudy (anonymous):

Ok, lets move onto the three questions now..

jimthompson5910 (jim_thompson5910):

how far did you get with #3 ?

OpenStudy (anonymous):

ill say sometimes also on question about linear

jimthompson5910 (jim_thompson5910):

do you agree that the original equation in #3 is a function? why or why not?

OpenStudy (anonymous):

Ok Wait, are we talking about the one all the way at the bottom or the one at the top?

OpenStudy (anonymous):

If we're talking about the question at the bottom, I say YES it is a function.

jimthompson5910 (jim_thompson5910):

the one with celsius and fahrenheit

jimthompson5910 (jim_thompson5910):

I'm thinking that's #3 since the first two were already done in another attachment

OpenStudy (anonymous):

Ok I see, we're basically on #1

jimthompson5910 (jim_thompson5910):

`If we're talking about the question at the bottom, I say YES it is a function.` correct. The one with sqrt(x-4) is a function. But they aren't asking if it's a function or not for this one

jimthompson5910 (jim_thompson5910):

yeah #1 if we're only focused on that attachment and forgetting everything else

OpenStudy (anonymous):

Yeah

jimthompson5910 (jim_thompson5910):

do you agree that the original equation in #1 is a function? why or why not?

OpenStudy (anonymous):

Yes, Its a linear function.

jimthompson5910 (jim_thompson5910):

why is it a linear function?

OpenStudy (anonymous):

I found the degree of the equation lol

jimthompson5910 (jim_thompson5910):

ok, another method is graphing and you'll notice a straight line forms or you can notice it's in the form y = mx+b where m is the slope, b is the y intercept m = 9/5 b = 32

jimthompson5910 (jim_thompson5910):

which ever method you use, the end result is that the original function is linear. How about the inverse? Is the inverse a function?

OpenStudy (anonymous):

Im not too sure on that part

jimthompson5910 (jim_thompson5910):

isolate C to find the inverse

OpenStudy (anonymous):

the inverse is not

jimthompson5910 (jim_thompson5910):

why not?

OpenStudy (anonymous):

I just took a guess just now, wasn't really sure..

jimthompson5910 (jim_thompson5910):

what do you get when you solve for C?

OpenStudy (anonymous):

\[- \frac{ 160 }{ 9 } + \frac{ 5F }{ 9 }\]

jimthompson5910 (jim_thompson5910):

notice how it's in the form y = mx+b now m = 5/9 b = -160/9

jimthompson5910 (jim_thompson5910):

so the inverse is linear and it's also a function

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

Yeah I see the formula you used. Lets move on?

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

okay #2

OpenStudy (anonymous):

My teacher was going over this but this specific problem is tricky

jimthompson5910 (jim_thompson5910):

what kind of function is f(x) ?

OpenStudy (anonymous):

a linear equation

jimthompson5910 (jim_thompson5910):

yep, and this rule is always true if you have a nonconstant linear function, then it always has an inverse function

jimthompson5910 (jim_thompson5910):

another rule \[\Large (f \circ f^{-1})(x) = x\] if the inverse function exists

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

so this means \[\Large (f \circ f^{-1})(-10) = -10\] because the inverse does exist

OpenStudy (anonymous):

most def! ok ok I see, lets move onto the last one..

jimthompson5910 (jim_thompson5910):

what is the domain of f(x)?

jimthompson5910 (jim_thompson5910):

hint: you cannot take the square root of a negative number

OpenStudy (anonymous):

Domain:

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

Rule: the domain of the original function will become the range of the inverse function this is because x and y swap

jimthompson5910 (jim_thompson5910):

so if the domain of f(x) is x >= 4, then the range of the inverse is x >= 4

OpenStudy (anonymous):

Most Def

OpenStudy (anonymous):

Thats what the range would be^

jimthompson5910 (jim_thompson5910):

yes the range of f(x) is y >= 0 so the domain of the inverse is x >= 0

jimthompson5910 (jim_thompson5910):

btw I meant to say `so if the domain of f(x) is x >= 4, then the range of the inverse is y >= 4` I changed the second `x>=4` to `y >= 4`

OpenStudy (anonymous):

thanks for the help.